Consider the general demand function: Qd=8000-16P +0.75M +30PR

Derive the equation for the demand function when M =$30,000 and PR=$50.


1
Expert's answer
2021-01-19T10:54:30-0500

Deriving the equation for the direct demand function when

M = $30,000 and PR = $50.


"Qd =8,000-16P + 0.75(30,000) + 30(50) = Qd =8,000-16P +24,000.00"


"Qd= 8,000-16P +0.75(30,000) + (22,500.00) +1500"



"Q_d =" $32,000.00-16p